The reaction 2H2S(g)⇌2H2(g)+S2(g), Kc=1.67×10−7, at 800∘C is carried out with the following initial concentrations: [H2S] = 0.375 M , [H2] =0.125 M , and [S2] = 0.000 M. Find the equilibrium [S2].

Express your answer with the appropriate units.
Answer #1

first construct the ICE table and substitute the values and then calculate the values as shown below
          2 H2S -----------------------------------> 2 H2      +    S2

I        0.375                                                0.125           0

C          -2X                                                   +2X            +X

E      (0.375-2X)                                          (0.125+2X)   X

Kc = [H2]2 [S2] / [H2S]2

1.67*10-7 = [0.125+2X]2 [X] / [0.375-2X]2

1.67*10-7*[0.375-2X]2 = [0.125+2X]2 [X]

=> x = 1.50290*10-6 M

answer => at equilibrium the [S2] is 1.50290*10-6 M